### v1

# A nonequilibrium quantum Otto engine enhanced via multi-parameter control

Abstract
Advances in experimental control of interacting quantum many-body systems with multiple tunable parameters—such as ultracold atomic gases and trapped ions—are driving rapid progress in quantum thermodynamics and enabling the design of quantum thermal machines. In this work, we utilize a sudden quench approximation as a means to investigate the operation of a quantum thermodynamic Otto cycle in which multiple parameters are simultaneously controllable. The method applies universally to many-body systems where such control is available, and therefore provides general principles for investigating their operation as a working medium in quantum thermal machines. We investigate application of this multi-parameter quench protocol in an experimentally realistic one-dimensional Bose gas as the working fluid, with control over both the frequency of an external harmonic trap and the interparticle interaction strength. We derive a general inequality for the net work of this two-parameter Otto cycle, demonstrating that this protocol out-performs its constituent single-parameter Otto cycles when operating as an engine, and additionally implying an enhancement to the coefficient of performance when operating as a refrigerator. Further, we demonstrate that multi-parameter control can exhibit dramatically improved performance of the Otto engine when compared not only to single-parameter constituent quenches but also to the combined effect of its constituent engine cycles.
I Introduction
Out-of-equilibrium dynamics of interacting many-body systems are at the forefront of both theoretical and experimental quantum physics, spurred in large part by rapid advancement in the experimental control over quantum platforms [1, 2, 3, 4, 5, 6, 7]. Investigation into such systems offers unprecedented opportunities within the field of quantum thermodynamics, which itself is a rapidly developing field with an emphasis on understanding thermodynamic principles arising within a quantum context [8, 9, 10, 11, 12, 13, 14]. Indeed, the realization of, and precise control over, many-body interacting quantum systems represents an important step in the study and understanding of thermodynamics in complex quantum systems.
Engine cycles have been central to the historical development of quantum thermodynamics, with origins dating back to 1953 in Scovil and Schulz-Dubois’ analysis of the maser as a single-body 3-level quantum heat engine [15]. Recently, in order to advance the understanding of quantum engines and their scaling, many researchers have focused on the operation of quantum devices in quantum many-body systems [16, 17, 6, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Such many-body interacting quantum systems enable the investigation of the role of uniquely quantum effects in engine operation, such as quantum coherence [30, 31, 32, 33, 34, 35] or correlations [36, 37, 38, 39, 19].
Notable recent examples have been the experimental realizations of quantum Otto engines in interacting ultracold Bose gases [16, 17]. These experiments exploited quasistatic control over both interparticle interaction strength and the frequency of an external harmonic trap in an alternating fashion, thus realizing uniquely quantum many-body thermodynamic devices. In particular, Ref. [16] elegantly demonstrated the role that quantum statistics played in the performance of their quasistatic quantum Otto engine cycle. Extending the investigation of quantum many-body engine cycles to non-quasistatic protocols, i.e. to out-of-equilibrium regimes of operation, is essential in order to understand their performance in scenarios that generate a finite power output, which vanishes in the quasistatic limit. However, such an extension remains challenging due to the typical complexity of simulating the out-of-equilibrium dynamics of interacting many-body systems in experimentally realistic parameter regimes, where precise and rapid control is available over various system parameters [40, 41, 42, 43, 28, 44, 20]. An extreme version of an out-of-equilibrium engine operation is realized via a sudden quench, where one can approximate the final post-quench state as unchanged from its initial thermal equilibrium state in order to again rely on exact thermal equilibrium expectation values for calculation of the net work and efficiency [19, 21].
In this work, we examine a quantum Otto engine cycle, as well as other related thermal machines, such as an Otto refrigerator, operating under a sudden quench of multiple externally tunable parameters. In particular, we extend the recent work on Otto engine cycles, operating under a single parameter quench for arbitrary quantum models in Ref. [21], to the case of simultaneous quenching of multiple control parameters. We apply this formalism to a harmonically trapped one-dimensional (1D) Bose gas, where we explore a simultaneous sudden quench of the harmonic trapping frequency and the interatomic interaction strength. This experimentally realizable engine cycle is shown to possess a region of dramatically enhanced performance compared not only to when quenching only a single parameter, but also to the combined performance under two single-parameter quenches.
II Multi-parameter sudden quench Otto engine
We begin by describing a general physical model that may correspond to, e.g., an ultracold quantum gas or a spin chain on a lattice, which incorporates multiple externally controllable parameters. In detail, in the context of second quantized quantum mechanics, we define a model Hamiltonian that consists first of the operators $\{\hat{\mathcal{V}}^{(\alpha)}\}$ (indexed by $\alpha$) intended to correspond to all operators with external control over a scalar strength parameter, denoted $\{c^{(\alpha)}\}$, such that the Hamiltonian contains the terms $c^{(\alpha)}\hat{\mathcal{V}}^{(\alpha)}$. The second set of operators, collectively denoted $\hat{H}_{0}$, corresponds to any and all terms without externally controlled strength parameters, a common example of which is the kinetic energy. Thus, the full Hamiltonian is given by,
$\hat{H}=\sum_{\alpha}c^{(\alpha)}\,\hat{\mathcal{V}}^{(\alpha)}+\hat{H}_{0}.$ (1)
We further note that the set $\{\hat{\mathcal{V}}^{(\alpha)}\}$ may consist of any combination of one-body operators, two-body operators, etc., and likewise any product of spin operators for spin models describing quantum magnetism.
The sudden quench approximation utilized in this work consists of assuming that the period of time over which the set of parameters $\{c^{(\alpha)}\}$ is quenched between the initial ($i$) and final ($f$) values, e.g. $\{c^{(\alpha)}_{i}\}\!\to\!\{c^{(\alpha)}_{f}\}$, is significantly faster than the timescale over which the density matrix of the initial equilibrium state, $\hat{\rho}_{i}$, is capable of responding. Under such an approximation, the energy of the state immediately after the quench may be evaluated as
$\langle\hat{H}\rangle_{f}\simeq\langle\hat{H}_{0}\rangle_{i}+\sum_{\alpha}c^{(\alpha)}_{f}\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{i},$ (2)
where $\langle\hat{\mathcal{O}}\rangle_{i}\!=\!\mathrm{Tr}[\hat{\rho}_{i}\hat{\mathcal{O}}$] denotes the expectation value of the operator $\hat{\mathcal{O}}$ in the equilibrium state defined by the initial density matrix $\hat{\rho}_{i}$. The work of the corresponding unitary stroke in the Otto engine cycle (see below), $W_{i\to f}\!\equiv\!\langle\hat{H}\rangle_{f}\!-\!\langle\hat{H}\rangle_{i}$, may therefore be approximated as
$W_{i\to f}\simeq\sum_{\alpha}(c^{(\alpha)}_{f}-c^{(\alpha)}_{i})\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{i},$ (3)
where the expectation value of the uncontrolled terms is unchanged, $\langle\hat{H}_{0}\rangle_{f}\!=\!\langle\hat{H}_{0}\rangle_{i}$, and therefore cancels when taking the difference $\langle\hat{H}\rangle_{f}\!-\!\langle\hat{H}\rangle_{i}$ in calculating the work. The work $W_{i\to f}$ is therefore determined entirely from the knowledge of expectation values $\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{i}$ calculated for the initial equilibrium state $\hat{\rho}_{i}$. We note that this method was recently utilized in Refs. [19, 21] to evaluate the performance of an Otto engine cycle under a sudden quench of the interparticle interaction strength, in which case the net work was shown to be proportional to the same-position particle-particle pair correlation function.
The multi-parameter quantum Otto engine cycle (described in greater detail in Appendix A) operates between high ($h$) and low ($l$) energy equilibrium states, facilitated by coupling of the working fluid to two external reservoirs. Generally, the type of coupling can be either thermal or diffusive, or both. Such an engine cycle consists of four strokes: two unitary work strokes in which the working fluid is isolated from all external reservoirs and the externally controlled parameters are suddenly quenched between $\{c_{l}^{(\alpha)}\}\rightleftarrows\{c_{h}^{(\alpha)}\}$; and two equilibration strokes where the working fluid is coupled with one of the two external reservoirs and the externally controlled parameters, $\{c^{(\alpha)}\}$, are kept constant at their post-quench values.
The net work, $W$, achieved during this engine cycle is calculated from the work input and output generated during the course of the two unitary work strokes described above, i.e., $W\!=\!W_{\mathbf{A}\to\mathbf{B}}\!+\!W_{\mathbf{C}\to\mathbf{D}}$. As each unitary work stroke is evaluated using the sudden quench approximation, given by Eq. (3), the total net work becomes:
$W\simeq-\sum_{\alpha}(c_{h}^{(\alpha)}-c_{l}^{(\alpha)})\left(\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}-\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{l}\right).$ (4)
Such an Otto engine cycle generates net beneficial work (done by the fluid) when W<0W\!<\!0. One may additionally evaluate the generalized engine efficiency, $\eta\!=\!-W/E_{\mathbf{B}\to\mathbf{C}}$ [45], using the calculated energy intake $E_{\mathbf{B}\to\mathbf{C}}$ during the equilibration stroke with the high energy reservoir, which is equal to $E_{\mathbf{B}\to\mathbf{C}}\!=\!\langle\hat{H}\rangle_{h}-(\,\langle\hat{H}\rangle_{l}+W_{\mathbf{A}\to\mathbf{B}}\,)$. This generalized efficiency is then given by
$\begin{split}\begin{aligned} \eta\simeq 1-\frac{\langle\hat{H}\rangle_{\mathit{h}}-\langle\hat{H}\rangle_{\mathit{l}}-\sum_{\alpha}(c^{(\alpha)}_{h}-c^{(\alpha)}_{l})\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{h}}{\langle\hat{H}\rangle_{h}-\langle\hat{H}\rangle_{l}-\sum_{\alpha}(c^{(\alpha)}_{h}-c^{(\alpha)}_{l})\big\langle\hat{\mathcal{V}}^{(\alpha)}\big\rangle_{l}},\end{aligned}\end{split}$ (5)
where we used the conservation of energy W+E=0W\!+\!E\!=\!0, with $E=E_{\mathbf{B}\to\mathbf{C}}\!+\!E_{\mathbf{D}\to\mathbf{A}}$ being the total energy. Importantly, the energy exchanged with the reservoirs during the equilibration strokes may take any form (e.g. heat, chemical work, etc.), requiring only that the set of strength parameters, $\{c^{(\alpha)}\}$, are kept constant. The case where energy intake and output is given entirely by heat corresponds to a conventional quantum Otto heat engine cycle, in which case the generalized efficiency is equivalent to the standard thermodynamic efficiency [45, 46]. Non-standard, or generalized, efficiency is utilized in cycles where the system-reservoir contact contains non-thermal components. A simple example of non-thermal contact is diffusive contact, which may be achieved through an exchange of particles between the system and reservoirs during the equilibration strokes; such a device is typically referred to as a thermo-chemical Otto engine [47, 48, 20, 21].
In the following we explore a particular case of this multi-parameter sudden quench Otto cycle in an experimentally realizable harmonically trapped 1D Bose gas. Such a system is amenable to rapid experimental control over both the frequency of the external harmonic trap and over the strength of interparticle interactions [7, 49, 50, 51, 26]. However, we emphasize here that the formalism outlined in this section applies universally to other quantum models capable of rapid control over multiple external parameters [21].
III The 1D Bose gas
The Lieb-Liniger model of a 1D Bose gas is a paradigmatic quantum model describing an ultracold atomic gas with contact pairwise interactions, being both experimentally realizable and having exact theoretical results [52, 53, 54, 7, 55, 56, 49, 54, 57, 58]. This model, in the uniform limit, consists first of an uncontrolled Hamiltonian operator, $\hat{H}_{0}$, corresponding to the kinetic energy,
$\begin{split}\begin{aligned} \hat{H}_{0}=-\frac{\hbar^{2}}{2m}\int dx\hat{\Psi}^{\dagger}(x)\frac{\partial^{2}}{\partial x^{2}}\hat{\Psi}(x).\end{aligned}\end{split}$ (6)
where $m$ is the atomic mass, and $\hat{\Psi}^{(\dagger)}(x)$ are the bosonic field annihilation (creation) operator at position $x$. Contact ($\delta$-function) interactions are incorporated in Eq. (1) via the interaction term
$c^{(2)}\hat{\mathcal{V}}^{(2)}=c^{(2)}\hat{\overline{G_{2}}}=c^{(2)}\!\int dx\hat{G}_{2}(x),$ (7)
where the two-body operator $\hat{\mathcal{V}}^{(2)}\equiv\hat{\overline{G_{2}}}=\int dx\hat{G}_{2}(x)$ corresponds to the integrated local (same-point) pair correlation $\hat{G}_{2}(x)\equiv\hat{G}_{2}(x,x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x)$, which we note does not depend on $x$ in a uniform system due to the translational invariance. The interaction strength $c^{(2)}$ is given by $c^{(2)}\!\equiv\!g/2\simeq\!\hbar\omega_{\perp}a_{s}$ away from confinement induced resonances [59] and is experimentally controllable via the tight transverse trapping frequency, $\omega_{\perp}$, or magnetic Feshbach resonance [60] tuning of the 3D $s$-wave scattering length $a_{s}$. We may utilize the normalized local (same-point) two-body correlation function,
$g^{(2)}(x)$ $=\frac{\langle\hat{G}_{2}(x)\rangle}{\rho(x)^{2}}=\frac{\langle\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x)\rangle}{\rho(x)^{2}},$ (8)
to express the expectation value of the integrated correlation function as $\langle\hat{\overline{G_{2}}}\rangle\!=\!\int dx\,g^{(2)}(x)\rho(x)^{2}$, where $\rho(x)=\langle\hat{\rho}(x)\rangle=\langle\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\rangle$ is the average particle number density. In the following, we utilize this to analytically evaluate the integrated local pair correlation function, taking advantage of known expressions for both the equilibrium density profile, $\rho(x)$ [61, 62], and for the normalized two-body correlation function, $g^{(2)}(x)$, in the relevant parameter regimes [63, 64, 65, 66, 67].
The uniform 1D Bose gas, with density $\rho\!=\!N/L$, at finite temperature is characterized by a dimensionless interaction strength, $\gamma\!=\!mg/\hbar^{2}\rho$, and a dimensionless temperature, $\tau\!=\!T/T_{d}$, where $T_{d}\!=\!\hbar^{2}\rho^{2}/2mk_{B}$ is the temperature of quantum degeneracy. This model becomes analytically tractable in six distinct regimes as a function of $\gamma$ and $\tau$ [65, 66, 67]. More generally, it is numerically solvable at all interaction strengths and temperatures in the thermodynamic limit through the thermodynamic Bethe ansatz (TBA) [53]. We note here that the interaction-driven quantum Otto cycle for this model in its uniform limit has been previously investigated under both adiabatic [26] and sudden quench [19, 21] protocols.
Experimental realization of the 1D Bose gas typically occurs within an external harmonic trapping potential [54, 7, 57], which is expressed via the one-body operator in Eq. (1),
$c^{(1)}\hat{\mathcal{V}}^{(1)}\equiv\hat{V}=\frac{1}{2}m\omega^{2}\int dx\,x^{2}\hat{\rho}(x),$ (9)
with $\hat{\mathcal{V}}^{(1)}\!\equiv\!\frac{1}{2}m\int dx\,x^{2}\hat{\rho}(x)$, where $c^{(1)}\!\equiv\!\omega^{2}$ corresponds to the harmonic trapping frequency squared and $\hat{\rho}(x)\!=\!\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)$ is the particle number density operator. We additionally define the atomic position variance, $\langle x^{2}\rangle\!\equiv\!\int dx\,x^{2}\langle\hat{\rho}(x)\rangle$, such that $\langle\hat{\mathcal{V}}^{(1)}\rangle\!=\!m\omega^{2}\langle x^{2}\rangle/2$.
In the presence of an external trapping potential, one may again utilize the numerically exact TBA introduced above, along with the additional assumption of a local density approximation [66], to accurately model the non-uniform system at finite temperature. Through the external harmonic trap, the density profile gains a dependence on the position, $\rho\!=\!\rho(x)$. The gas may then be described in terms of the dimensionless interaction strength and temperature at the trap center, i.e. $\gamma_{0}\!=\!mg/\hbar^{2}\rho(0)$ and $\tau_{0}\!=\!2mk_{B}T/\hbar^{2}\rho(0)^{2}$, and the average total number of particles in the system, $N\!=\!\int dx\langle\hat{\rho}(x)\rangle$ [66]. The interaction-driven and separately the volumetric (i.e. harmonic frequency quench) quantum Otto cycles for this system have been previously investigated under both quasi-static and sudden quench protocols [20, 21, 27, 17, 16].
Here, we instead consider a scenario with rapid control over both the interaction strength and harmonic trapping frequency between two fixed values, denoted $(g_{l},g_{h})$ and $(\omega^{2}_{l},\omega^{2}_{h})$, with $g_{l}\!\leq\!g_{h}$ and $\omega^{2}_{l}\!\leq\!\omega^{2}_{h}$. This enables realization of the two-parameter sudden quench quantum Otto engine cycle. The net work of such an engine cycle is calculated via Eq. (4) (see Appendix A), giving
$-W\simeq$ $\frac{1}{2}(g_{h}-g_{l})\left(\langle\hat{\overline{G_{2}}}\rangle_{h}-\langle\hat{\overline{G_{2}}}\rangle_{l}\right)$
$+\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\left(\langle x^{2}\rangle_{h}-\langle x^{2}\rangle_{l}\right).$ (10)
Though the cycles are inherently interconnected, with the thermal equilibrium expectation values $\langle\hat{\overline{G_{2}}}\rangle_{h(l)}$ and $\langle x^{2}\rangle_{h(l)}$ depending on both controllable parameters $g$ and $\omega$, we separate this formula into an interaction-driven sub-cycle, $W^{g}\!=\!\frac{1}{2}(g_{h}-g_{l})\big(\langle\hat{\overline{G_{2}}}\rangle_{h}-\langle\hat{\overline{G_{2}}}\rangle_{l}\big)$, and a volumetric sub-cycle, $W^{\omega}=\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\left(\langle x^{2}\rangle_{h}-\langle x^{2}\rangle_{l}\right)$, such that $W\!\simeq\!-(W^{g}+W^{\omega})$. Such a partition will be useful later for examining the enhancement of engine performance when considering a simultaneous quench of both controllable parameters relative to single-parameter quenches.
The net work can therefore be deduced either in the ground state system, where the temperatures of the external reservoirs, and therefore the working fluid, are fixed to $T=0$, and only particle exchange takes place between the working fluid and the reservoirs, or at finite temperatures via TBA calculation, where both heat and particles may be exchanged. In both cases, one may calculate the equilibrium expectation values of both $\langle\hat{\overline{G_{2}}}\rangle$ and $\langle x^{2}\rangle$ in the high- and low-energy equilibrium states. Further, there is a corresponding result for the generalized efficiency, given by Eq. (5), again calculable via the TBA. We note that interaction enhancement of a volumetric Otto engine cycle was previously studied in the context of an adiabatic quantum Otto cycle for few-body interacting systems in Ref. [44].
IV Two-parameter enhancement
Before turning to calculation of the net work, we examine the properties of this two-parameter Otto cycle in comparison with the two single-parameter Otto cycles, taken separately. These single-parameter quantum Otto cycles correspond to those commonly considered in thermodynamics [21], consisting of control over only the ratio $g_{h}/g_{l}$ (or $\omega_{h}^{2}/\omega_{l}^{2}$) while keeping the complementary quantity constant, i.e., $\omega_{h}^{2}/\omega_{l}^{2}\!=\!1$ (or $g_{h}/g_{l}\!=\!1$). Through this, we will demonstrate that the net work of the two-parameter Otto cycle is always greater than the sum of the single-parameter cycles taken in isolation. We then generalize the arguments presented for the harmonically trapped 1D Bose gas to general quantum models describing scenarios where the two-parameter cycle out-performs the sum of the single-parameter in terms of net work.
For clarity, we emphasize the dependence of the equilibrium states $(h)$ and $(l)$ on the interaction strength and harmonic trapping frequency as $\langle\cdot\rangle_{(g,\omega)}$. To illustrate this notation, we show the total system energy, $\langle\hat{H}\rangle_{(g,\omega)}$, in Fig. 1, and additionally demonstrate the path in parameter space which both the single-parameter and two-parameter Otto cycles traverse. Using this notation, the work input ($W_{\mathbf{A}\to\mathbf{B}}$) of the interaction-driven single-parameter Otto cycle is
$W_{\mathbf{A}\to\mathbf{B}}^{g}=$ $\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{l},\omega_{l})}.$ (11)
Similarly, the work input of the volumetric single-parameter Otto cycle is
$W_{\mathbf{A}\to\mathbf{B}}^{\omega}=$ $\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{l})}.$ (12)
Note that we have chosen to begin all cycles in the same low energy equilibrium state, characterized by the same temperature, particle number, and set of parameters $g_{l}$ and $\omega_{l}$.
Figure 1: Total system energy ($\langle\hat{H}\rangle_{(g,\omega)}$) diagram for the sudden quench Otto cycle with a harmonically trapped 1D Bose gas as the working fluid. Here, control is over both the interaction strength, $g$, and the external harmonic trapping frequency, $\omega$, given on the horizontal and vertical axes, respectively, for realistic experimental values [68]. Beginning in the low energy equilibrium state, given by $(g_{l},\omega_{l})$, one may consider a single-parameter Otto cycle where either interaction strength is suddenly quenched, $(g_{l},\omega_{l})\!\to\!(g_{h},\omega_{l})$, corresponding to an interaction-driven Otto cycle, or the harmonic trapping frequency is quenched, $(g_{l},\omega_{l})\!\to\!(g_{l},\omega_{h})$, corresponding to a volumetric Otto cycle. In contrast, the two-parameter Otto cycle consists of quenching both parameters, $(g_{l},\omega_{l})\!\to\!(g_{h},\omega_{h})$, resulting in an enhanced performance (see text) over the sum of the single-parameter Otto cycles.
We are interested in the difference in net work between the two-parameter Otto cycles and the sum of the single-parameter Otto cycles, which we denote
$\Delta W=(-W)-(-W^{g}-W^{\omega}).$ (13)
For the case of engine operation, having $\Delta W\!>\!0$ implies that the net work generated by two-parameter Otto cycle is greater than the sum of the single-parameter cycles taken together.
Inspecting first the work input from the compression stroke, $W_{\mathbf{A}\to\mathbf{B}}$, of the two-parameter Otto cycle,
$W_{\mathbf{A}\to\mathbf{B}}\simeq$ $\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{l},\omega_{l})}$
$+\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{l})},$ (14)
we observe that this is exactly the sum of the work inputs of the interaction-driven and volumetric single-parameter Otto cycles, given in Eqs. (11) and (12), respectively. This results from the fact that, regardless of cycle type, the system begins in the same low energy equilibrium state and implies that $\Delta W>0$ cannot originate from the compression stroke.
Therefore, a nonzero difference in net work $\Delta W$, if any, can only originate from the difference in the work outputs in the expansion stroke, $W_{\mathbf{C}\to\mathbf{D}}$, of these Otto engine cycles,
$-\Delta W=(-W_{\mathbf{C}\to\mathbf{D}})-(-W_{\mathbf{C}\to\mathbf{D}}^{g}-W_{\mathbf{C}\to\mathbf{D}}^{\omega}).$ (15)
For the two-parameter Otto cycle, the work output in the expansion stroke is given by
$-W_{\mathbf{C}\to\mathbf{D}}$ $=\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}$
$\quad+\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{h},\omega_{h})}.$ (16)
In contrast, work outputs of the single-parameter Otto cycles, which keep the complementary parameters constant, are given by
$W_{\mathbf{C}\to\mathbf{D}}^{g}=\frac{1}{2}(g_{h}-g_{l})\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})},$ (17)
for the interaction-driven single-parameter Otto cycle, and by
$W_{\mathbf{C}\to\mathbf{D}}^{\omega}=\frac{1}{2}m(\omega_{h}^{2}-\omega_{l}^{2})\langle x^{2}\rangle_{(g_{l},\omega_{h})}.$ (18)
for the volumetric single-parameter Otto cycle. Substituting Eqs. (IV), (17), and (18) into Eq. (15) for $\Delta W$, gives:
$-\Delta W$ $=\frac{1}{2}(g_{h}-g_{l})\left(\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}-\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}\right)$ (19)
$+\frac{1}{2}(\omega_{h}^{2}-\omega_{l}^{2})\left(\langle x^{2}\rangle_{(g_{h},\omega_{h})}-\langle x^{2}\rangle_{(g_{l},\omega_{h})}\right).$ (20)
This expression enables a direct comparison between the net work of the single-parameter Otto cycles with the sub-cycles of the two-parameter Otto cycle, $W^{g}$ and $W^{\omega}$, where $W\!=\!W^{g}\!+\!W^{\omega}$, as defined after Eq. (III).
To proceed with such a comparison, we note that an increase to the harmonic trapping frequency, $\omega_{l}\!\to\!\omega_{h}$ with $\omega_{h}\!>\!\omega_{l}$, for fixed interaction strength $g_{h}$, compresses the atomic cloud, thereby increasing the integral of the squared density profile, $\int dx\,\rho(x)^{2}$. To guarantee an overall increase to the integrated correlation function, given below Eq. (8) as $\langle\hat{\overline{G_{2}}}\rangle\!=\!\int dx\,g^{(2)}(x)\rho(x)^{2}$, we additionally require that $g^{(2)}(x)$ grows under this same compression. Indeed, we know that $g^{(2)}(x)$ monotonically decreases with the dimensionless interaction strength $\gamma(x)\!=\!mg/\hbar^{2}\rho(x)$ [67]. As this dimensionless interaction strength is inversely proportional to the density, we find that the total integrated correlation function indeed increases with the trapping frequency, meaning $\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{h})}\!>\!\langle\hat{\overline{G_{2}}}\rangle_{(g_{h},\omega_{l})}$ for $\omega_{h}\!>\!\omega_{l}$. Therefore, the net work extracted from the interaction-driven sub-cycle of the two-parameter Otto cycle, given on the first line of Eq. (IV), exceeds that extracted from the interaction-driven single-parameter cycle in Eq. (17).
Similarly, increasing the interaction strength, $g_{l}\!\to\!g_{h}$ with $g_{h}\!>\!g_{l}$, at a fixed value of the harmonic trapping frequency, $\omega_{h}$, broadens the atomic cloud, as increased interparticle repulsion drives the atoms further apart. This results in an increased second moment of the density distribution, $\langle x^{2}\rangle$, meaning $\langle x^{2}\rangle_{(g_{h},\omega_{h})}\!>\!\langle x^{2}\rangle_{(g_{l},\omega_{h})}$. From this, we find that the contribution to $W_{\mathbf{C}\to\mathbf{D}}$ from the volumetric sub-cycle of the two-parameter Otto cycle, given on the second line of Eq. (IV), exceeds $W_{\mathbf{C}\to\mathbf{D}}^{\omega}$ for the single-parameter Otto cycle, given in Eq. (18).
Therefore, by comparing both terms contributing to the two-parameter Otto cycle against their single-parameter counterparts, we arrive at the conclusion that $W_{\mathbf{C}\to\mathbf{D}}\!>\!-W_{\mathbf{C}\to\mathbf{D}}^{g}\!-\!W_{\mathbf{C}\to\mathbf{D}}^{\omega}$, meaning $\Delta W\!>\!0$. Thus, for the case of a sudden quench of both interaction strength and harmonic trapping frequency in a 1D Bose gas, the net work of the two-parameter quantum Otto cycles is greater than the sum of that for the single-parameter Otto cycles taken in isolation. We additionally note that the same arguments apply to any Otto cycle where control is over both interaction strength and volume of the gas. In particular, the arguments presented above also apply to the the 1D Bose gas confined to either a uniform box trap or within a ring, where the total system length is controllable. Such a scenario may be thought of as a quantum analogue of the classical piston engine.
More generally, one may consider whether there is a benefit to operating any two-parameter sudden quench Otto cycle over the related single-parameter cycles. In particular, in Appendix B, we examine an Otto cycle where the sudden quench is taken for the scalar strength parameters $c^{(1)}$ and $c^{(2)}$, as first introduced for a general physical model in Eq. (1), and which are associated with the operators $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$, respectively. The net work of this sudden quench Otto cycle is given by Eq. (4), for a sudden quench over the parameters $c^{(1)}$ and $c^{(2)}$ between their respective values in the low ($l$) and high ($h$) energy equilibrium states. Following arguments similar to those laid out above for the case of a 1D Bose gas, we find that the net work of the two-parameter Otto cycle out-performs the sum of the single-parameter Otto cycles if both $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$ grow under an increase to their complementary strength parameters, i.e. $c^{(2)}$ and $c^{(1)}$, respectively.
An application for such an inequality, beyond the 1D Bose gas already discussed, is the harmonically trapped 1D Fermi gas, commonly known as the Yang-Gaudin model [69, 70]. There, one may additionally utilize control over the spin polarization to enhance performance, as recently explored for a uniform system in Ref. [21]. Further, one may consider applying a two-parameter quench in the context of the transverse-field Ising model, where it is possible to exploit control over both the nearest-neighbor interaction and the transverse magnetic field [71, 72, 73]. While application to these models would be of interest, in the following we restrict our attention to the case of a harmonically trapped 1D Bose gas.
Finally, we note that the enhanced performance of a two-parameter quench thermal machine over its single-parameter constituents, expressed as $\Delta W>0$, remains true regardless of whether the Otto Cycle operates as an engine (which is what we are considering here, in the main text), refrigerator, thermal accelerator, or heater [19], (see Appendix B for further details, where we illustrate this for an Otto refrigerator).
V Zero temperature quasicondensate regime
To gain analytical insight and obtain transparent quantitative results on the enhancement of the Otto engine performance under a two-parameter quench, we first consider a simple example of a harmonically trapped, zero-temperature 1D Bose gas in the weakly interacting regime, with the dimensionless interaction strength in the trap centre satisfying $\gamma_{0}\ll 1$) [74, 75, 66].
Since the working fluid is in its zero temperature ground state, the Otto cycle under consideration cannot operate as a heat engine (which cycles between cold and hot thermal equilibrium states via exchange of heat with the cold and hot reservoirs). Instead, we examine its operation in the chemical engine scenario [76, 77, 47, 48, 20], facilitated by chemical work due to exchange of particles when the working fluid is in contact with the reservoirs for attaining the low- ($l$) and high-energy ($h$) equilibrium states.
At zero temperature, the weakly interacting Bose gas may is described by a Thomas-Fermi approximation [62, 61]. From this, one is able to calculate both equilibrium expectation values, $\langle x^{2}\rangle$ and $\langle\hat{\overline{G_{2}}}\rangle$, present in Eq. (III) analytically. More specifically, for the expectation value $\langle x^{2}\rangle$ we obtain $\langle x^{2}\rangle\!=\!(N/5)R_{\mathrm{TF}}^{2}$, where $R_{\mathrm{TF}}\!=\!(3Ng/2m\omega^{2})^{1/3}$ is the Thomas-Fermi radius [61] (for further details, see Appendix C). For the correlation function $\langle\hat{\overline{G_{2}}}\rangle$, on the other hand, we obtain the approximation $\langle\hat{\overline{G_{2}}}\rangle\simeq\!g^{(2)}(0)\int dx\rho(x)^{2}\!\simeq\!bN\rho(0)g^{(2)}(0)$, where $\rho(0)\!=\!(9mN^{2}\omega^{2}/32g)^{1/3}$ is the density of the Thomas-Fermi profile at the trap center, and $b$ is a constant factor determined by the form of the density profile, with b=4/5b\!=\!4/5 for the Thomas-Fermi inverted parabola. Thus, the integrated correlation function depends on
$g^{(2)}(0)$ $=\frac{\langle\hat{\Psi}^{\dagger}(0)\hat{\Psi}^{\dagger}(0)\hat{\Psi}(0)\hat{\Psi}(0\rangle}{\rho(0)^{2}},$ (21)
which is the normalized local two-body correlation function at the trap center, i.e. at x=0x\!=\!0. In the weakly interacting ground state, we may approximate this correlation function as $g_{l(h)}^{(2)}(0)\!\simeq\!1\!-\!2\sqrt{\gamma_{0,l(h)}}/\pi$ [67], where $\gamma_{0,l(h)}\!=\!mg_{l(h)}/\hbar^{2}\rho_{l(h)}(0)$ is the dimensionless interaction strength at the trap center in the low (high) energy equilibrium state, which depends on the peak density of the equilibrium density profile, $\rho_{l(h)}(x)$.
Finally, to also evaluate engine efficiency via Eq. (5), we require an expression for the total energy, $\langle\hat{H}\rangle_{l(h)}\!=\!\langle\hat{V}\rangle_{l(h)}+(g_{l(h)}/2)\langle\hat{\overline{G_{2}}}\rangle_{l(h)}$, in the low (high) energy equilibrium states. This can be expressed in the Thomas-Fermi approximation as
$\langle\hat{H}\rangle_{l(h)}\!=\!E^{\mathrm{TF}}_{l(h)}\!+\!\frac{g_{l(h)}}{2}bN_{l(h)}\rho_{l(h)}(0)\left(g^{(2)}_{l(h)}(0)\!-\!1\right),$ (22)
where we have defined $E^{\mathrm{TF}}_{l(h)}\!=\!3N_{l(h)}g_{l(h)}\rho_{l(h)}(0)/5$, as the total energy in the low (high) energy equilibrium state of the harmonically trapped 1D Bose gas that is fully coherent, i.e. $g^{(2)}_{l(h)}(0)\!=\!1$ [61, 62]. The second term contains corrections to the interaction energy due to the approximate expression for the correlation function, where $g^{(2)}_{l(h)}(0)\!-\!1\!\simeq\!-2\sqrt{\gamma_{0,l(h)}}/\pi$.
Figure 2: Performance of the two-parameter sudden quench quantum Otto engine cycle in the weakly interacting ground state of the 1D Bose gas using the Thomas-Fermi approximation. Panels (a) and (b) demonstrate the net work ($W$) and efficiency ($\eta$), respectively, as a function of both the interaction strength ratio, $g_{h}/g_{l}$, along the horizontal axis, and harmonic trapping frequency ratio, $\omega_{h}^{2}/\omega_{l}^{2}$, along the vertical axis. The net work is represented in harmonic oscillator units defined by the longitudinal frequency in the low energy equilibrium state, $\omega_{l}$. The low energy equilibrium state ($l$) is parameterized by N=2000N\!=\!2000 total particles at dimensionless interaction strength $\gamma_{0}\!=\!4.9\times 10^{-2}$. $\Delta N\!=\!200$ particles are exchanged with the reservoirs while in contact.
Combining these approximations in Eqs. (5) and (III), we thus obtain simple analytic results for the efficiency and net work of the two-parameter sudden quench chemical Otto engine and illustrate them in Fig. 2, as functions of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, the system is initialized in the low-energy equilibrium state with N=2000N\!=\!2000 particles. Then, following the work input stroke, $W_{\mathbf{A}\to\mathbf{B}}$, the working fluid takes in $\Delta N\!=\!200$ particles during the equilibration stroke $\mathbf{B}\!\to\!\mathbf{C}$ with the high energy reservoir. To operate in a closed cycle, the same number of particles is later output into the low energy reservoir during the corresponding equilibration stroke $\mathbf{D}\!\to\!\mathbf{A}$ [78].
The net work of this two-parameter sudden quench Otto cycle, shown in Fig. 2(a), indicates that engine operation occurs only within a finite region of the quenched parameter ratios, with a maximum net work achieved at the center of this region. In particular, the maximum net work achieved is $W/\hbar\omega_{l}\!\simeq\!4.9\times 10^{4}$, which, when normalized to the total particle number in the low energy equilibrium state, corresponds to $W/N_{l}\hbar\omega_{l}\!\simeq\!25$. Further, we find an efficiency at maximum net work of $\eta\!\simeq\!0.42$. Normalized to total particle number in the low energy equilibrium state, Importantly, this performance represents a significant enhancement over the single-parameter sudden quench quantum Otto engine cycles investigated recently for the 1D Bose gas in Ref. [21].
Figure 3: Comparison of the two-parameter sudden quench quantum Otto engine cycle against the single-parameter Otto engine cycles. In panel (a), we demonstrate the net work of the single-parameter Otto engine cycles as a function of their quenched parameter; the interaction-driven cycle, $W_{g}$, is shown as the solid blue line, whereas the volumetric cycle, $W_{\omega}$, as the solid red line. The value of the quenched parameter ratio is denoted ‘ratio’. This is contrasted with the maximum net work of the two-parameter Otto engine cycle, denoted $W^{\mathrm{max}}$ and shown as the solid yellow line. The maximum net work of both single-parameter Otto cycles are given by the blue and red dashed lines for the interaction-driven and volumetric cycles, respectively, with their sum shown as the black dashed line. In contrast, the maximum net work of the two-parameter Otto engine cycle is greater than this sum by more than an order of magnitude. Panel (b) demonstrates the efficiency of the single-parameter Otto engine cycles, with colors corresponding to those shown in panel (a). The efficiency at maximum net work of the two-parameter Otto cycle is shown as the solid yellow line, and clearly out-performs both single-parameter cycles both in terms of magnitude and breadth of operation.
In particular, the net work and efficiency under control of only the ratio $g_{h}/g_{l}$ (or $\omega_{h}^{2}/\omega_{l}^{2}$) while keeping the complementary quantity, $\omega_{h}^{2}/\omega_{l}^{2}\!=\!1$ (or $g_{h}/g_{l}\!=\!1$), constant, are shown in Fig. 3 (a) and (b). The quenched parameter is denoted as the ‘ratio’ on the horizontal axis of this figure, and denotes either $g_{h}/g_{l}$ for $W_{g}$, or $\omega_{h}^{2}/\omega_{l}^{2}$ for $W_{\omega}$. For comparison, we plot the maximum net work of the two-parameter Otto engine cycle, $W^{\mathrm{max}}$, which is found by calculating the volumetric ratio $\omega_{h}^{2}/\omega_{l}^{2}$ that gives the maximum net work for each value of the interaction strength ratio $g_{h}/g_{l}$.
Previously, in Sec. IV, we have shown that, for a working fluid consisting of a harmonically trapped 1D Bose gas, the performance of the two-parameter Otto cycle generally out-performs the sum of the individual cycles taken in isolation. Here, in this concrete example of a two-parameter Otto cycle, we see that this engine greatly outperforms the individual Otto cycles, as the two-parameter cycle generates net work that is more than an order of magnitude greater than the sum of both individual engine cycles operating at their respective maxima (shown as the black dashed line in Fig. 3(a)). Additionally, though it was not guaranteed by the analysis presented in Sec. IV, we find that the efficiency of the two-parameter Otto engine cycle also outperforms what is achieved by both single-parameter engine cycles.
To explain why engine operation only occurs over a region of the parameter space, as shown in Fig. 2 (a), we decompose the net work into two contributions,
$W\simeq W_{\mathrm{coh.}}+W_{\mathrm{corr.}},$ (23)
where $W_{\mathrm{coh.}}$ denotes the extractable net work from a fully coherent gas, i.e. approximating $g^{(2)}_{l}(0)\!=\!g^{(2)}_{h}(0)\!\simeq\!1$. In particular, utilizing the analytic expressions derived from the Thomas-Fermi approximation, the net work extracted from a fully coherent working fluid is given by
$W_{\mathrm{coh.}}=$ $-\frac{b}{2}(g_{h}-g_{l})\left(N_{h}\rho_{h}(0)-N_{l}\rho_{l}(0)\right)$
$-\frac{1}{10}m(\omega_{h}^{2}-\omega_{l}^{2})\left(N_{h}R_{h}^{2}-N_{l}R_{l}^{2}\right).$ (24)
The second term in Eq. (23) contains higher order terms arising from corrections to the local second order correlation function due to the finite interaction strength,
$W_{\mathrm{corr.}}\!=\!\frac{b}{\pi}(g_{h}\!-\!g_{l})\left(N_{h}\rho_{h}(0)\sqrt{\gamma_{0,h}}\!-\!N_{l}\rho_{l}(0)\sqrt{\gamma_{0,l}}\right).$ (25)
We now analyze these two contributions in greater detail.
To simplify our analysis, we consider the net work as a function of equal parameter ratios, that is for $g_{h}/g_{l}\!=\!\omega_{h}^{2}/\omega_{l}^{2}\!\equiv\!r$. Applying this to the coherent contribution to the net work, we find
$W_{\mathrm{coh.}}$ $\!\simeq\!-E^{\mathrm{TF}}_{l}(r\!-\!1)\left(\frac{N_{h}^{5/3}}{N_{l}^{5/3}}\!-\!1\right).$ (26)
Therefore, assuming that we are operating with a finite quench size, r>0r\!>\!0, and with non-zero particle intake, $\Delta N=N_{h}\!-\!N_{l}>0$, the coherent contribution to the net work guarantees engine operation, W<0W\!<\!0.
The additional terms contributing to the net work, given in Eq. (25), contain terms arising from the corrections to the coherent correlations, i.e. from $g^{(2)}(0)\!-\!1\!=\!2\sqrt{\gamma(0)}/\pi$. It is these corrections to the correlation function that restrict engine operation to a limited range of the parameter ratios. In particular, the dimensionless interaction strength of the high energy equilibrium state, denoted $\gamma_{0,h}$, scales as $\gamma_{0,h}\!\propto\!r$, causing this contribution to reduce the overall net work. At large enough values of this ratio, these correction terms result in W>0W\!>\!0, meaning the cycle no longer operates as an engine. We therefore note that taking account of these corrections to coherent correlations is essential to correctly evaluating the performance of this quantum Otto cycle.
VI Finite temperature quasicondensate
At low but finite temperatures such that $2\gamma_{0}\!\ll\!\tau_{0}\!\ll\!2\sqrt{\gamma_{0}}$, the weakly interacting ($\gamma_{0}\!\ll\!1$) 1D Bose gas inhabits the thermal quasicondensate regime [67]. Here, the Thomas-Fermi approximation for the density profile is no longer a good approximation, and must be replaced by the thermodynamic Bethe ansatz solution under a local density approximation [53, 66]. This method is numerically exact, and enables calculation of both equilibrium expectation values required to evaluate engine performance via Eqs. (5) and (III). At non-zero temperature, we may additionally utilize a finite temperature difference between the two equilibrium states $\mathbf{A}$ and $\mathbf{C}$ (see Appendix A), meaning our working fluid exchanges both heat and particles with the reservoirs during equilibration, i.e. we investigate operation of a thermo-chemical Otto engine cycle [20].
Figure 4: Performance of the two-parameter sudden quench thermo-chemical quantum Otto engine cycle in the finite temperature quasicondensate regime, evaluated via numerically exact TBA methods. Panels (a) and (b) demonstrate the net work and efficiency, respectively, as a function of the ratio of both quenched parameters. The parameter values for the low energy equilibrium state are chosen to match those used in Fig. 2, but at a finite dimensionless temperature of $\tau_{0}\!=\!1.2\!\times\!10^{-2}$ (see text). Further, we utilize a fixed temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$, making this a thermo-chemical quantum Otto cycle. We observe a maximum net work of $W/\hbar\omega_{l}\!\simeq\!1.1\!\times\!10^{4}$, corresponding to $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$, and an efficiency at maximum net work of $\eta\!\simeq\!0.04$. Panels (c) and (d) display a comparison between the maximum net work and efficiency at maximum net work of the two parameter Otto cycle, respectively, against the single-parameter Otto cycles, as previously shown for the zero temperature quasicondensate system in Fig. 3.
Performance of the two-parameter thermo-chemical quantum Otto cycle, in terms of net work and efficiency, for a harmonically trapped 1D Bose gas in the quasicondensate regime is shown in Fig. 4 (a) and (b). Here, we use the same total net particle exchange of $\Delta N\!=\!200$ as in Fig. 2, but with an additional temperature ratio between the high and low energy equilibrium states of $T_{h}/T_{l}\!=\!1.33$. We note that, in Ref. [19], it was found that—for a harmonically trapped system (as opposed to the uniform case)—it was essential to enable diffusive contact and chemical work in order that net beneficial work could be extracted from the interaction-driven Otto cycle under a sudden quench.
We observe that, when compared to the pure chemical quantum Otto cycle, operating at $T=0$ and shown in Fig. 2, the two-parameter thermo-chemical quantum Otto cycle investigated here operates as an engine for a broader range of both quenched parameter ratios. This difference stems from the fact that we are incorporating a finite temperature difference between the high and low energy equilibrium states, in addition to the particle number difference. This increase to temperature in the high energy equilibrium state broadens the atomic density profile, increasing the second moment of the density distribution $\langle x^{2}\rangle$, and hence resulting in an enhancement to the net work extracted from the volumetric sub-cycle (see Eq. (III)).
The interaction-driven sub-cycle is similarly altered when operating at finite temperature. However, the changes arising in this sub-cycle due to finite-temperature operation represent a smaller alteration to the net work and efficiency compared with the modifications in the volumetric sub-cycle, as finite-temperature corrections to the local second-order correlation function are minimal in the finite-temperature quasicondensate regime [67]. We additionally note that, upon reducing the temperature ratio between the high and low energy equilibrium states to the point where $T_{h}/T_{l}\!=\!1$, we would effectively reproduce the results shown for the zero temperature quasicondensate system, investigated in Fig. 2, but including corrections to the correlation function and density profile arising from finite temperature effects [66, 67].
In panels (c) and (d) of Fig. 4, we contrast the maximum work and efficiency at maximum work of the two-parameter Otto cycle against the performance of the single-parameter Otto cycles in the finite-temperature quasicondensate regime. Through this, we observe a significant enhancement to the performance when utilizing a two-parameter quench, similar to what was observed in zero-temperature quasicondensate system in Fig. 3. Notably, the maximum value of net work achieved from this two-parameter Otto engine cycle is again approximately an order of magnitude greater than the sum of that achieved by the single-parameter Otto engine cycles, shown as the black dashed line in Fig.  4  (c), with a similar improvement to the efficiency at maximum work.
Though the improved performance is guaranteed for all parameter regimes of the 1D Bose gas, it is not a priori clear whether the enhancement achieved remains as significant outside of the weakly interacting quasicondensate regime. We therefore turn to examining engine operation, utilizing the same numerically exact TBA methods, in the strongly interacting Tonks-Girardeau regime [79, 80, 67].
Figure 5: Performance of the two-parameter sudden quench Otto engine cycle for a harmonically trapped 1D Bose gas in the strongly interacting Tonks-Girardeau regime, calculated via numerically exact TBA methods. All panels are presented in the same form as in Fig. 4. The low energy equilibrium state is fixed by $N_{l}\!=\!20$, $\tau_{0}\!\simeq\!0.18$, and $\gamma_{0}\!\simeq\!8.5$. The high energy equilibrium state has $N_{h}\!=\!22$, and $T_{h}/T_{l}\!=\!2$. We observe a maximum net work of $W/\hbar\omega_{l}\!\simeq\!1.1$, corresponding to $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5\times 10^{-2}$, and an efficiency at maximum net work of $\eta\!\simeq\!0.02$.
VII Tonks-Girardeau gas
The strongly interacting ($\gamma_{0}\gg 1$), low temperature ($\tau_{0}\!\ll\!\pi^{2}/(1+2/\gamma_{0})^{2}$) regime [67], commonly known as the Tonks-Girardeau gas [79, 80], is well approximated by a nearly ideal Fermi gas due to the strong repulsion between bosons [80, 65, 81]. Experimental realization of the Tonks-Girardeau gas requires extremely strong transverse confinement, and is typically realized in the form of an array of 1D Bose gases within a 2D optical lattice at low total atom numbers [82, 49, 55, 83, 7]. Yet, despite the low total number of atoms, the system remains well described by the TBA under the same local density approximation [66, 49], which we again utilize to examine operation of the two-parameter Otto engine.
Engine performance in the Tonks-Girardeau regime is demonstrated in Fig. 5, again in terms of net work and efficiency as a function of both quenched parameter ratios, $g_{h}/g_{l}$ and $\omega_{h}^{2}/\omega_{l}^{2}$. Here, our working fluid consists of N=20N\!=\!20 total atoms, at dimensionless interaction strength $\gamma_{0}\!\simeq\!8.5$ and dimensionless temperature $\tau_{0}\!=\!0.18$, meaning the system inhabits the low temperature regime of fermionization [67]. The Otto cycle investigated here is again thermo-chemical, where $\Delta N\!=\!2$ particles are exchanged with the reservoirs during the equilibration strokes, whereas the high and low equilibrium states are chosen to have a temperature ratio of $T_{h}/T_{l}\!=\!2$.
Notably, the enhancement to net work and efficiency achieved via the two-parameter cycle in the Tonks-Girardeau gas is not as significant when contrasted to operation in the quasicondensate regime. To compare the two-parameter engine cycle between these regimes, we consider the maximum net work normalized to the total particle number in the low energy equilibrium state, $N_{l}$, which varies significantly between the Tonks-Girardeau ($N_{l}\!=\!20$) and quasicondensate ($N_{l}\!=\!2000$) working fluids. Upon doing this, we find that the Otto cycle operating in the Tonks-Girardeau regime achieves a maximum net work of $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5\!\times\!10^{-2}$, which is two orders of magnitude less than the same cycle operating in the quasicondensate regime, where $W/N_{l}\hbar\omega_{l}\!\simeq\!5.5$ (see Fig. 4 caption).
The reduction in performance in the Tonks-Girardeau regime may be attributed to the effect of fermionization of the interaction-driven sub-cycle. In particular, under strong interparticle interactions, the local second-order correlation function, and therefore the interaction energy of the working fluid, is reduced to near zero, dramatically reducing the net work of the interaction-driven sub-cycle. Further, adding particles to a fermionic system is associated with a large intake of energy, in contrast to adding particles to a bosonic system, where most particles condense into low energy states. Hence, in the Tonks-Girardeau regime, particle intake from the high energy reservoir is associated with a large energy penalty, and results in low efficiencies.
VIII Conclusions
In this work, we investigated the operation of a sudden quench Otto cycle with control over multiple external parameters. To do this, we extended the work done recently in Ref. [21], for sudden quenches of single parameters, to the general case of arbitrary sets of controllable parameters. Under the sudden quench approximation, the total net work separates into a sum of its constituent quenches. From this, we derived general principles dictating when the net work extracted from a two-parameter sudden quench protocol exceeds that from related single-parameter cycles taken in isolation for arbitrary quantum models, benefiting an Otto engine performance. Further, this enhancement to the net work of the two-parameter Otto cycle was found to result in an improved coefficient of performance when operating as an Otto refrigerator.
The methods introduced were applied to the case of an Otto engine cycle in an experimentally realizable harmonically trapped 1D Bose gas with contact interactions. Control over both the strength of interactions and the harmonic trapping frequency was shown to result in a region of enhanced performance when both parameters are simultaneously quenched. This was analytically illustrated in the Thomas-Fermi approximation in the simplest example of a chemical engine operating in the weakly interacting regime of the 1D Bose gas at zero temperature.
The same enhancement to both net work and efficiency was demonstrated numerically using the thermodynamic Bethe ansatz in a finite temperature quasicondensate and the strongly interacting Tonks-Girardeau regime of the 1D Bose gas operating as a thermochemical engine. Finally, we highlight that the methods introduced here are applicable universally to quantum mechanical models, relying only on rapid control over multiple external parameters.
Acknowledgments
The authors thank Lewis Williamson and Matthew Davis for valuable discussions. This work was supported through Australian Research Council Discovery Project Grant Nos. DP190101515 and DP240101033.
Appendix A Sudden quench Otto cycle
The multi-parameter quantum Otto cycle, operating between two reservoirs denoted $(h)$ (higher energy) and $(l)$ (lower energy) in Fig. 6, consists of the following four strokes,
  • (1)
    Unitary compression, $\mathbf{A}\!\to\!\mathbf{B}$: the working fluid, which is initially in an equilibrium state $\hat{\rho}_{l}$ at a set of Hamiltonian strength parameters $\{c_{l}^{(\alpha)}\}$, is disconnected from the reservoir $(l)$ and has its strength parameters suddenly quenched from $\{c_{l}^{(\alpha)}\}\!\to\!\{c_{h}^{(\alpha)}\}$, with $c_{h}^{(\alpha)}\!>\!c_{l}^{(\alpha)}$ for all $\alpha$, and energy difference $\langle\hat{H}\rangle_{\mathbf{B}}\!-\!\langle\hat{H}\rangle_{\mathbf{A}}\!>\!0$. This means that the work $W_{\mathbf{A}\to\mathbf{B}}\!=\!\langle\hat{H}\rangle_{\textbf{B}}\!-\!\langle\hat{H}\rangle_{\textbf{A}}\!>\!0$ is done on the fluid. Here, $\langle\hat{H}\rangle_{\textbf{J}}$ is the expectation value of the total Hamiltonian given by Eq. (1), i.e., the total internal energy of the system, in state $\textbf{J}=\{\textbf{A,B,C,D}\}$ shown in Fig. 6.
  • (2)
    Thermalization with reservoir (h), $\mathbf{B}\!\to\!\mathbf{C}$: the working fluid, now in an out-of-equilibrium state, is connected to reservoir (h) and is allowed to equilibrate while keeping the strengths, $\{c^{(\alpha)}_{h}\}$, constant. The working fluid takes in energy $E_{\mathbf{B}\to\mathbf{C}}\!=\!\langle\hat{H}\rangle_{\textbf{C}}\!-\!\langle\hat{H}\rangle_{\textbf{B}}\!>\!0$ from the reservoir.
  • (3)
    Unitary expansion, $\mathbf{C}\!\to\!\mathbf{D}$: working fluid, now in an equilibrium state described by $\hat{\rho}_{h}$, is decoupled from reservoir (h) and has its strength parameters suddenly quenched $\{c^{(\alpha)}_{h}\}\!\to\!\{c^{(\alpha)}_{l}\}$, resulting in work $W_{\mathbf{C}\to\mathbf{D}}\!=\!\langle\hat{H}\rangle_{\textbf{D}}\!-\!\langle\hat{H}\rangle_{\textbf{C}}\!<\!0$ done by the fluid.
  • (4)
    Thermalization with reservoir (l), $\mathbf{D}\!\to\!\mathbf{A}$: the nonequilibrium working fluid is connected to reservoir (l), allowing for energy exchange at constant $\{c^{(\alpha)}_{l}\}$, thus ejecting energy $E_{\mathbf{D}\to\mathbf{A}}\!=\!\langle\hat{H}\rangle_{\textbf{A}}\!-\!\langle\hat{H}\rangle_{\textbf{D}}\!<\!0$ into the reservoir, and returning to its original equilibrium state $\hat{\rho}_{l}$.
Figure 6: Internal energy of the working fluid, $\langle\hat{H}\rangle$, in an interaction-driven quantum many-body Otto engine cycle operating between two interaction strengths $c_{l}$ and $c_{h}$ and in cyclic connection ($\mathbf{B}\!\to\!\mathbf{C}$ and $\mathbf{D}\!\to\!\mathbf{A}$) to two reservoirs denoted (l) for the low energy and (h) for the high energy state of the working fluid. Unitary work strokes $\mathbf{A}\!\to\!\mathbf{B}$ and $\mathbf{C}\!\to\!\mathbf{D}$ are denoted via dashed lines to signify the fact that these strokes are accomplished via a sudden quench rather than by passing through the intermediate equilibrium states tracing these lines.
The net work of this sudden quench Otto engine cycle is given by Eq. (4) in the main text. Such a cycle generates net beneficial work (done by the fluid) when W<0W\!<\!0, with a generalised engine efficiency given by Eq. (5). Notably, this generalized efficiency accounts for the fact that the energy exchange with the reservoirs may take any form (e.g. heat, chemical work, etc.).
Appendix B General multi-parameter enhancement for engine and refrigerator operation
Here, we examine an Otto cycle where the sudden quench is taken for the scalar strength parameters $c^{(1)}$ and $c^{(2)}$, as first introduced for a general physical model in Sec. II, and which are associated with the operators $\hat{\mathcal{V}}^{(1)}$ and $\hat{\mathcal{V}}^{(2)}$, respectively. The net work of this sudden quench Otto cycle is given by Eq. (4) of the main text, for a sudden quench over the parameters $c^{(1)}$ and $c^{(2)}$ between their respective values in the low ($l$) and high ($h$) energy equilibrium states.
To examine whether the net work of the two-parameter cycle exceeds the sum of the single-parameter cycles, we follow the same arguments laid out in the main text. In particular, we compare the net work of the two-parameter sub-cycle associated with $c^{(1)}$ with the single-parameter cycle associated with quenching only parameter $c^{(1)}$. The difference in net work for these contributions is given by
$-\Delta W^{(1)}\!\simeq$ $(c_{h}^{(1)}-c_{l}^{(1)})$
$\times$ $\left(\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{h}\right)}\!-\!\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{l}\right)}\right)$, (27)
which must be greater than zero for the two-parameter cycle to out-perform the single-parameter protocol. In particular, we require that $\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{h}\right)}\!>\!\big\langle\hat{\mathcal{V}}^{(1)}\big\rangle_{\left(c^{(1)}_{h},c^{(2)}_{l}\right)}$, i.e. we require the equilibrium expectation value of $\hat{\mathcal{V}}^{(1)}$ to increase with the complementary parameter, $c^{(2)}$. If this condition is met for both sub-cycles, then we arrive at the inequality
$\Delta W=-\Delta W^{(1)}-\Delta W^{(2)}>0,$ (28)
for the net work extracted from two-parameter Otto cycle to exceed the sum of the single-parameter cycles, where $\Delta W$ is defined in Eq. (13).
Finally, we note that, if the inequality $\Delta W\!>\!0$ is valid for a particular quantum model operating in an Otto cycle, it remains valid regardless of whether the Otto cycle operates as an engine, or one of the alternative protocols (e.g., refrigerator, thermal accelerator, or heater [19]). Indeed, considering the refrigerator operation as one of the alternative protocols [45, 19], shown schematically in Fig. 7, we first note that the key parameter is the coefficient of performance ($\mathrm{CoP}$), given by
$\mathrm{CoP}_{\mathrm{two-param.}}=\frac{E_{\mathbf{D}\to\mathbf{A}}}{W}.$ (29)
Inspecting the combination of the two single-parameter Otto cycles, we may assign a coefficient of performance to the operation of both single-parameter cycles connected to the same high and low energy reservoirs. This combined coefficient of performance is therefore given by
$\mathrm{CoP}_{\mathrm{one-param.}}\!=\!\frac{E_{\mathbf{D}\to\mathbf{A}}^{g}+E_{\mathbf{D}\to\mathbf{A}}^{\omega}}{W^{g}+W^{\omega}}.$ (30)
Assuming that the net work of the two-parameter Otto cycle exceeds the sum of the single-parameter cycles, i.e. assuming $\Delta W\!=\!-W-(-W^{g}+-W^{h})\!>\!0$, and recalling that W+E=0W\!+\!E=0 for cyclic operation, where $E\!=\!E_{\mathbf{B}\to\mathbf{C}}\!=\!E_{\mathbf{D}\to\mathbf{A}}$ (see Sec. II), upon subtracting the net work of the single-parameter cycles from that of the two-parameter cycle, we find that $E\!-(\!E^{g}\!+\!E^{\omega})>0$.
Figure 7: Refrigerator operation for the two-parameter Otto cycle. For this protocol, the Otto cycle is the same as in Fig. 6 for the engine, except that now the energy flow between the working fluid and reservoirs is reversed [45], meaning equilibration with the low energy reservoir takes in energy, i.e. $E_{\mathbf{D}\to\mathbf{A}}\!>\!0$, implying that the reservoir ($l$), that thermalizes the working fluid to its low-energy thermal equilibrium state, is being cooled down by the working fluid.
Refrigerator operation assumes that we have a finite energy intake from the high energy reservoir, as energy must be extracted for cooling to occur [45]. This means that the energy intake $E_{\mathbf{B}\to\mathbf{C}}$ must be positive. We may then subtract this positive contribution from the inequality $E\!-(\!E^{g}\!+\!E^{\omega})>0$ to arrive at
$E_{\mathbf{D}\to\mathbf{A}}\!-\!(E_{\mathbf{D}\to\mathbf{A}}^{g}\!\!+\!\!E_{\mathbf{D}\to\mathbf{A}}^{\omega})\!>\!0.$ (31)
Hence, the numerator of the two-parameter coefficient of performance, given in Eq. (29), exceeds that of the one-parameter coefficient of performance in Eq. (30).
Likewise, as we are assuming that the working fluid satisfies the inequality $\Delta W\!=\!-W-(-W^{g}+-W^{h})\!>\!0$, the denominator of the two-parameter coefficient of performance, Eq. (29), is strictly less than that of the one-parameter case, Eq. (30). Hence, we find that, for the case where net work of the two-parameter Otto refrigerator cycle exceeds that of the sum of the single-parameter refrigerator cycles,
$\mathrm{CoP}_{\mathrm{two-param.}}\!>\!\mathrm{CoP}_{\mathrm{one-param.}}.$ (32)
meaning that refrigerator operation, like engine operation, is enhanced for the two-parameter protocol over the combined effects of single-parameter operation.
Appendix C Thomas-Fermi approximation
Here, we derive the various analytic formulas employed in the main text for the investigation of a quantum Otto engine cycle in the quasicondensate regime. In particular, we utilize the Thomas-Fermi approximation for the density profile [62, 61],
$\rho(x)=\rho(0)\left(1-\frac{x^{2}}{R_{\mathrm{TF}}^{2}}\right),$ (33)
where $\rho(0)\!=\!(9mN^{2}\omega^{2}/32g)^{1/3}$ is the peak density at the trap center (i.e. $x=0$), and $R_{\mathrm{TF}}\!=\!(3Ng/2m\omega^{2})^{1/3}$ is the 1D Thomas-Fermi radius. This approximation is valid for the ground state density profile in the weakly interacting quasicondensate regime, and remains a good approximation at finite but sufficiently low temperatures, i.e. for $\tau_{0}\!\ll\!2\gamma_{0}$ with $\gamma_{0}\!\ll\!1$ [67, 66, 62, 61].
The volumetric Otto engine cycle explored in the main text required evaluation of the atomic position variance, given by
$\langle x^{2}\rangle$ $=\int dx\rho(x)x^{2}.$ (34)
A straightforward calculation based on the Thomas-Fermi approximation introduced above gives
$\langle x^{2}\rangle=\frac{N}{5}R_{TF}^{2}.$ (35)
This formula was utilized recently to explore the volumetric Otto engine cycle under a sudden quench in Ref. [21].
To evaluate the performance of the interaction-driven Otto engine, we require the total correlation function, defined in the main text as
$\langle\hat{\overline{G_{2}}}\rangle=\int dx\,\langle\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x)\rangle.$ (36)
To progress, we introduce the normalized two-body correlation function, which is analytically tractable in the various asymptotic regimes that the 1D Bose gas possesses [65],
$g^{(2)}(x)$ $=\frac{\langle\hat{\Psi}^{\dagger}(x)\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)\hat{\Psi}(x\rangle}{\rho(x)^{2}}.$ (37)
Rearranging the above, we integrate to find an expression for the total integrated correlation function,
$\langle\hat{\overline{G_{2}}}\rangle=\int dx\,g^{(2)}(x)\rho(x)^{2}.$ (38)
We then utilize the fact that the normalized correlation function, $g^{(2)}(x)$, only slowly varies over the atomic density profile. As such, we may evaluate the total correlation as,
$\langle\hat{\overline{G_{2}}}\rangle\simeq g^{(2)}(0)\int dx\,\rho(x)^{2},$ (39)
which was first utilized in Ref. [65], and remains a good approximation for the total correlation over the entire parameter space of the 1D Bose gas.
Next, in the weakly interacting quasicondensate regime, we know that the correlation function is well approximated by the totally coherent value of $g^{(2)}(0)\simeq 1$. Indeed, though a more detailed expression is known [65, 67], this represents a higher-order correction, and does not provide additional insights to the simple explanation given in the main text.
Finally, we utilize the Thomas-Fermi density profile to evaluate the squared density profile in Eq. (39) as
$\int dx\,\rho(x)^{2}=bN\rho(0),$ (40)
where $b\equiv\,\int dx\,\rho(x)^{2}/N\rho(0)$ is a constant factor determined by the density profile, and is given by b=4/5b\!=\!4/5 in the quasicondensate regime. We then arrive at the final expression for the total integrated correlation function utilized in the main text,
$\langle\hat{\overline{G_{2}}}\rangle\simeq bN\rho(0).$ (41)